# Polarized multiwavelength emission from pulsar wind – accretion disk interaction in a transitional millisecond pulsar

### v1

@@ Appendix F Polarization Properties in the Low Mode; Constraints on Polarized Radio Emission in the two Modes
In the scenario outlined in Section 4.1.4, the upper limit on the radio linear polarization fraction in the high mode is consistent with partially self-absorbed synchrotron radiation from the optically thick portion of the compact jet. The exact cause of the switch to the low mode remains uncertain, though the detection of short millimeter-band flares during a few of these switches in past observations points to sudden ejections of magnetized plasma, likely originating from the innermost regions of the accretion flow. The ejection of this matter could thus force the shock region to move outward, reducing the X-ray flux and the pulsed emission. As the plasma propagates downstream in the compact jet, it becomes optically thin in the radio band [10]. This framework also accounts for the optical polarization properties measured during the low mode. If the shock moves outward, it may still contribute a measurable amount of polarized light. Moreover, the optical polarization angle observed in the low mode is consistent with that of the high mode within uncertainties, suggesting a common underlying emission mechanism for the optical linear polarization across both modes and reinforcing the idea that residual shock emission remains active in the low mode. In the soft X-ray band, where the jet could dominate up to 100% of the emission [10], the maximum expected polarization would be 25%. The upper limit for the X-ray polarization degree during the low mode, % at a 90% c.l., makes it impossible to draw firm conclusions from this measurement. Alternatively, the switch to the low mode may result from matter occasionally penetrating the light cylinder, temporarily forcing the system into a regime where the rapidly rotating magnetic field anchored to the NS propels the inflowing matter away. As matter expands outward, it becomes increasingly rarefied and more transparent to synchrotron emission at radio wavelengths [122].
@@
At radio wavelengths, synchrotron radiation is expected to produce linear polarization fractions of up to 10% for optically thick emission and up to 70% for optically thin emission, assuming uniform magnetic fields and a power-law electron distribution with an index of 2 [85]. In accreting stellar-mass black holes, polarization fractions are typically below 10%, but can reach up to 50% during instances of optically thin emission ([68] and references therein). Thus, the upper limit on the linear polarization measured from J1023 in the high mode is consistent with the presence of an optically thick compact jet, similar to what is observed in other X-ray binaries. In contrast, the lower polarization limit during the low mode may result from partial cancellation of orthogonally polarized jet components due to misaligned magnetic fields.
Appendix G Polarimetric properties of accreting neutron stars in LMXBs
IXPE has observed several accreting NSs in LMXBs, revealing diverse polarimetric behaviors. The polarization degree of the soft X-ray emission typically remains around a few percent across all observed systems. In this section, we summarize the polarization properties of the observed systems and briefly discuss their interpretations.
G.1 Z-sources
Cyg X2.
Cyg X-2 exhibits stable spectral characteristics along the Z-track, with a notable hard X-ray excess above 30 keV during the Horizontal Branch phase linked to episodic radio jet emissions. IXPE measurements indicate a polarization degree of and a PA of , aligning with the radio jet direction. The Comptonized emission component shows a polarization degree of , also aligned with the jet, suggesting the primary polarization originates from a spreading layer on the NS surface, with potential contributions from disk reflection [48].
Sco X1.
Sco X1 is the brightest persistent extrasolar X-ray source, approaching the Eddington luminosity for a 1.4 NS. It has a 19-hour orbital period and it is located 2.13 kpc away. IXPE observations in the soft state revealed an average PD of and a PA of . The PD of the accretion disk emission was constrained to 3.2%, while that of the Comptonized component was [78].
GX 51.
GX 51 is located near the Galactic Center and is notable for radio emission from a compact jet. IXPE observations in the horizontal and normal-flaring branches of the Z-track measured PDs of and , respectively. A 20∘ variation in PA with energy was observed, likely due to differing polarization angles of the disk and Comptonization components peaking at various energies [46].
GX 3400.
GX 3400 is known for its stable spectral behavior and associated radio jet emissions. IXPE observations detected significant X-ray polarization when the source was mainly within the horizontal branch, with a PD of and a PA of in the 2–8 keV range. A variation in polarization angle with energy was measured, which was attributed to either the accretion disk and Comptonized emissions, or a mix of blackbody emission and Comptonized emission [13]. Alternatively, the low-energy polarization could be due to the accretion disk whereas the high energy polarization could arise from the corona [79].
Subsequent observations during the normal branch phase recorded a PD of and a PA of , with the PD exhibiting a marginal increase with energy. The PA is consistent with that measured in the horizontal branch. The observed polarized emission is consistent with originating from blackbody emission, Comptonized emission, or reflection features, with no significant contribution from the accretion disk [13].
G.2 Atoll sources
GS 1826238.
GS 1826238 remained in a hard spectral state until its transition to a high soft state in early 2016. IXPE observations in 2022 set a upper limit on the PD of 1.3% [25].
GX 31
GX 31 was observed with IXPE, along with NuSTAR and NICER, in August 2024. No significant X-ray polarization was detected in the 2–8 keV band, with an upper limit of 1.3% on the PD at the 99% c.l.. The X-ray spectra are well described by a model consisting of thermal emission from an accretion disk, a Comptonized emission component, and a component accounting for reflected photons off the disk. The upper limits on the polarized signal from the disk and the Comptonization component are consistent with theoretical expectations for a system with a standard disk and a Comptonizing region that likely has a geometry consistent with a spreading layer [53].
GX 99.
GX 99 exhibits a consistent bright soft state and a 4.2-hour modulation in both optical and X-ray bands. A spectro-polarimetric analysis of IXPE data detected X-ray polarization with a PD of 1.7%0.4% in the 2–8 keV band, which increased from a value consistent with null polarization in the 2–4 keV range to 3.2% in the 4–8 keV range. Spectral modeling indicated that the emission comprises thermal radiation from the accretion disk and a Comptonized component from an optically thick corona, suggesting that the polarization primarily originates from the Comptonized component [29]. A subsequent study measured a PD of 2.2%0.5% in the 4–8 keV band. Broadband spectral modeling required thermal disk emission, a Comptonized component, and a reflection component. An upper limit of 4% was found on the PD of the disk emission; both Comptonization near the NS and disk reflection were found to contribute to the observed polarization [118].
4U 162449.
4U 162449 is a persistent system known for its periodic 6–8 hour dips every 21 hours and flaring behavior above the bright persistent emission. IXPE detected an average PD of and a PA of during non-dip intervals, with PD increasing from (4–6 keV) to (6–8 keV). This suggests polarization arises from Comptonization in an extended, geometrically thin slab corona covering part of the accretion disk and disk reflection. During X-ray dips, an upper limit of 22% was established for the PD [108, 54].
G.3 Ultracompact binaries
4U 162667.
4U 162667 is a 7.66-s pulsar in an ultracompact binary with a 42-minute orbital period. IXPE observations constrained the pulse-averaged PD to 4% (95% confidence). However, phase-resolved spectro-polarimetric analysis revealed a marginal polarization detection in the power-law component with a PD of (90% confidence; [88]).
4U 1820303.
4U 1820303 is an ultra-compact low-mass X-ray binary located in the globular cluster NGC 6624 at 8 kpc. The system features a 11-minute orbital period and exhibits luminosity variations over a 170-day super-orbital cycle. Its spectrum comprises a blackbody or multicolor disk component, a Comptonization component, and an iron K line around 6.5 keV. IXPE observations measured a PD of with a PA of in the 4–7 keV band and with a PA of in the 7–8 keV band. spectro-polarimetric analysis suggests orthogonal polarization between the accretion disk and the hard spectral component from the boundary/spreading layer [40].
G.4 Transient systems
XTE J1701462.
This transient system experienced an outburst in September 2022, and was observed twice by IXPE with a 10-day interval. The initial observation was performed when the source was in the horizontal branch and recorded an average PD of , increasing from (2–4 keV) to (4–8 keV). The subsequent observation, taken in the normal branch, showed a reduced average PD of 0.6%. Spectral modeling incorporated an unpolarized thermal emission from the accretion disk and a Comptonized component with a PD of , indicating the Comptonized emission likely originates from a spreading or boundary layer above the NS surface, decreasing at later outburst stages [32, 72, 133]. A recent time-resolved polarimetric analysis revealed a stable PD of 2% in the second observation along with significant variations in the PA (yielding depolarization in the time-averaged data and hence the smaller average PD). These results point to rapid changes in the geometry of the corona, possibly from a slab geometry to a more vertically extended spreading layer geometry [134].
SRGA J144459.2604207.
SRGA J144459.2604207 is an accreting MSP detected in outburst in February 2024. Over a 10-day monitoring period, IXPE measured an average PD of in the 2–8 keV band, with a PA of . The polarization exhibited significant energy dependence, reaching between 3 and 6 keV, while remaining below 2% (90% c.l.) in the 2–3 keV range. No significant variability in the polarization properties was detected across spin and orbital phases. The observations are consistent with a moderately fan-beam emission originating from two localized hotspots with a small magnetic obliquity of . An upper limit on the PD of 8.5% (90% c.l.) was set during type I bursts [99].
G.5 Other systems
GX 131. GX 131 is a bright, persistent Galactic type I X-ray burster. IXPE observations in October 2023 detected a PD of (2–8 keV) with a PA of . A follow-up observation in February 2024 measured an overall PD of and a PA of . The polarization was attributed to a combination of Comptonization in the boundary or spreading layer and scattering from the accretion disk wind [17, 15].

### v2

@@ Appendix E Polarization Properties in the Low Mode; Constraints on Polarized Radio Emission in the two Modes
In the scenario outlined in Section 4.1.3, the upper limit on the radio linear polarization fraction in the high mode is consistent with partially self-absorbed synchrotron radiation from the optically thick portion of the compact jet. The exact cause of the switch to the low mode remains uncertain, though the detection of short mm-band flares during a few of these switches in past observations points to sudden ejections of magnetized plasma, likely originating from the innermost regions of the accretion flow. The ejection of this matter could thus force the boundary region to move outward, reducing the X-ray flux and the pulsed emission. As the plasma propagates downstream in the compact jet, it becomes optically thin in the radio band (M. C. Baglio et al., 2023). This framework also accounts for the optical polarization properties measured during the low mode. If the boundary region moves outward, it may still contribute a measurable amount of polarized light. Moreover, the optical polarization angle observed in the low mode is consistent with that of the high mode within uncertainties, suggesting a common underlying emission mechanism for the optical linear polarization across both modes and reinforcing the idea that residual emission from the collision remains active in the low mode. In the X-ray band, where the jet could dominate up to 100% of the emission (M. C. Baglio et al., 2023), the maximum expected polarization would be 25%. The upper limit for the X-ray polarization degree during the low mode, % at a 90% c.l., makes it impossible to draw firm conclusions from this measurement.
Alternatively, the switch to the low mode may result from matter occasionally penetrating the light cylinder, temporarily forcing the system into a regime where the rapidly rotating magnetic field anchored to the NS propels the inflowing matter away. As matter expands outward, it becomes increasingly rarefied and more transparent to synchrotron emission at radio wavelengths (A. Veledina et al., 2019).
@@
At radio wavelengths, synchrotron radiation is expected to produce linear polarization fractions of up to 10% for optically thick emission and up to 70% for optically thin emission, assuming uniform magnetic fields and a power-law electron distribution with an index of 2 (M. S. Longair, 2011). In accreting stellar-mass black holes, polarization fractions are typically below 10%, but can reach up to 50% during instances of optically thin emission (A. K. Hughes et al. 2023 and references therein).